# Doubling the size of quantum selected configuration interaction based on seniority-zero space and its application to QC-QSCI-AFQMC

### v1

@@ Acknowledgements
This work was supported by TOPPAN Digital Inc. This work was also supported by MEXT Quantum Leap Flagship Program (MEXTQLEAP) Grant No. JPMXS0120319794, the JST COI-NEXT Program Grant No. JPMJPF2014, the JST ASPIRE Program Grant No. JPMJAP2319, and the New Energy and Industrial Technology Development Organization (NEDO) Grant No. JPNP20017. This research was partially supported by the JSPS Grants-in-Aid for Scientific Research (KAKENHI) Grant No. JP23H03819. We thank the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo for the use of the facilities. This work was also conducted using SQUID at the Cybermedia Center, the University of Osaka.

### v2

@@ Acknowledgements
This work was supported by MEXT Quantum Leap Flagship Program (MEXTQLEAP) Grant No. JPMXS0120319794, JST COI-NEXT Program Grant No. JPMJPF2014, JST ASPIRE Program Grant No. JPMJAP2319, and New Energy and Industrial Technology Development Organization (NEDO) Grant No. JPNP20017. This research was partially supported by the JSPS Grants-in-Aid for Scientific Research (KAKENHI) Grant No. JP23H03819. We thank the Supercomputer Center, the Institute for Solid State Physics at the University of Tokyo for the use of the facilities. This work was also conducted using SQUID at the Cybermedia Center at the University of Osaka.